PLEASE I NEED HELP WITH THIS EQUATION X=2Y-1
Y-X=-5
AND I HAVEN'T GOT VERY FAR.

Answer by stanbon(48568)

X=2Y-1
Y-X=-5
--------
Rearrange:
x-2y=-1
x-y=5
--------
Subtract the 1st equation from the 2nd to get:
y=6
--------
Substitute that into either of the original equations to get:
x=2y-1
x=2*6-1
x=11
==========
solution x=11, y=6
-------
Check that solution in the OTHER original equation:
x-y=5
11-6=5
5=5
===========
Cheers,
Stan H.